can you find the next 3 term in this sequence
1,5,4,20,19,_,_,_

Answer by checkley71(8403)

THE SEQUENCE IS *5,-1
1*5=5
5-1=4
4*5=20
20-1=19
[19*5=95
95-1=94
94*5=470]